# Retraining with Predicted Hard Labels Provably Increases Model Accuracy

### v1

@@ Appendix B Proof of Theorem 1
$\mathbb{E}\Big{[}\exp\big{(}-r{\mathrm{y}}_{1}\langle{\mathbf{x}}_{1},{\mathbf{x}}\rangle\big{)}\Big{]}$ $\leq\exp\big{(}-\mu\sigma r{\mathrm{x}}^{(1)}\big{)}\exp\Bigg{(}\frac{(\sigma r\|{\mathbf{x}}\|)^{2}}{2}\Bigg{)}.$ (21)
Next, plugging in equations (20) and (21) into eq. 13, we get:
$\mathbb{E}\Big{[}\exp\big{(}r\widehat{\mathrm{y}}_{1}\langle{\mathbf{x}}_{1},{\mathbf{x}}\rangle\big{)}\Big{]}$ $\leq\exp\Bigg{(}\frac{(\sigma r\|{\mathbf{x}}\|)^{2}}{2}\Bigg{)}\Big{(}(1-p)\exp\big{(}\mu\sigma r{\mathrm{x}}^{(1)}\big{)}+p\exp\big{(}-\mu\sigma r{\mathrm{x}}^{(1)}\big{)}\Big{)}$ (22)
$=\exp\Bigg{(}\frac{(\sigma r\|{\mathbf{x}}\|)^{2}}{2}+\mu\sigma r{\mathrm{x}}^{(1)}\Bigg{)}\Big{(}1-p+p\exp\big{(}-2\mu\sigma r{\mathrm{x}}^{(1)}\big{)}\Big{)}$ (23)
$\leq\exp\Bigg{(}\frac{(\sigma r\|{\mathbf{x}}\|)^{2}}{2}+\mu\sigma r{\mathrm{x}}^{(1)}-p+p\exp\big{(}-2\mu\sigma r{\mathrm{x}}^{(1)}\big{)}\Bigg{)},$ (24)
where the last step follows because $\exp(z)\geq 1+z$ for all $z\in\mathbb{R}$. Substituting eq. 24 into eq. 12 followed by eq. 11 yields:
$\mathbb{E}\Big{[}\exp\big{(}r\varepsilon\big{)}\Big{]}\leq\exp\Bigg{(}n\Bigg{\{}\frac{(\sigma r\|{\mathbf{x}}\|)^{2}}{2}+\mu\sigma r{\mathrm{x}}^{(1)}-p+p\exp\big{(}-2\mu\sigma r{\mathrm{x}}^{(1)}\big{)}\Bigg{\}}\Bigg{)}.$ (25)
Plugging in eq. 25 into eq. 10, we get:
$\mathbb{P}\big{(}\varepsilon\leq 0\big{)}\leq\exp\Bigg{(}n\Bigg{\{}\underbrace{\frac{(\sigma r\|{\mathbf{x}}\|)^{2}}{2}+\mu\sigma r{\mathrm{x}}^{(1)}-p+p\exp\big{(}-2\mu\sigma r{\mathrm{x}}^{(1)}\big{)}}_{:=\nu}\Bigg{\}}\Bigg{)},$ (26)
for any $r<0$. Let $\nu:=\frac{(\sigma r\|{\mathbf{x}}\|)^{2}}{2}+\mu\sigma r{\mathrm{x}}^{(1)}-p+p\exp\big{(}-2\mu\sigma r{\mathrm{x}}^{(1)}\big{)}$. We will pick $r=-c\Big{(}\frac{\mu{\mathrm{x}}^{(1)}}{\sigma\|{\mathbf{x}}\|^{2}}\Big{)}$ for some $c\in(0,1)$. Putting this into the expression of $\nu$, we get:
$\nu=\mu^{2}\Bigg{(}\frac{{\mathrm{x}}^{(1)}}{\|{\mathbf{x}}\|}\Bigg{)}^{2}\Big{(}\frac{c^{2}}{2}-c\Big{)}+p\Bigg{(}\exp\Bigg{(}2c\mu^{2}\Bigg{(}\frac{{\mathrm{x}}^{(1)}}{\|{\mathbf{x}}\|}\Bigg{)}^{2}\Bigg{)}-1\Bigg{)}$ (27)
For brevity, let $u=\mu^{2}\Big{(}\frac{{\mathrm{x}}^{(1)}}{\|{\mathbf{x}}\|}\Big{)}^{2}$. Note that $u\leq\frac{1}{4}$, as $\mu\leq\frac{1}{2}$ and ${\mathrm{x}}^{(1)}\leq\|{\mathbf{x}}\|$. Rewriting the above equation in terms of $u$, we get:
$\nu=u\Big{(}\frac{c^{2}}{2}-c\Big{)}+p\big{(}\exp(2cu)-1\big{)}$ (28)
Note that $2cu\leq\frac{1}{2}$ as $c\leq 1$ and $u\leq\frac{1}{4}$. Therefore, applying Lemma 1 above, we get:
$\nu\leq u\Big{(}\frac{c^{2}}{2}-c\Big{)}+p\big{(}2cu+4c^{2}u^{2}\big{)}\leq\underbrace{u\Big{(}\frac{c^{2}}{2}-c\Big{)}+p\big{(}2cu+c^{2}u\big{)}}_{\text{(A)}},$ (29)
where the last step is obtained by using the fact that $u\leq\frac{1}{4}$. (A) is minimized by choosing $c=\frac{1-2p}{1+2p}$. With this choice of $c$, we get:
$\nu\leq-\frac{u}{2}\Bigg{(}\frac{(1-2p)^{2}}{1+2p}\Bigg{)}.$ (30)
Further using the fact $1+2p\leq 2$ and plugging in the value of $u=\mu^{2}\Big{(}\frac{{\mathrm{x}}^{(1)}}{\|{\mathbf{x}}\|}\Big{)}^{2}$ above, we get:
$\nu\leq-\frac{\mu^{2}}{4}(1-2p)^{2}\Bigg{(}\frac{{\mathrm{x}}^{(1)}}{\|{\mathbf{x}}\|}\Bigg{)}^{2}.$ (31)
Plugging this into eq. 26 and recalling that $\big{|}\textup{cos}\big{(}\beta({\mathbf{x}})\big{)}\big{|}=\frac{|{\mathrm{x}}^{(1)}|}{\|{\mathbf{x}}\|}$, we get:
$\mathbb{P}\big{(}\varepsilon\leq 0\big{)}\leq\exp\Bigg{(}-\frac{n\mu^{2}}{4}(1-2p)^{2}\textup{cos}^{2}\big{(}\beta({\mathbf{x}})\big{)}\Bigg{)}.$ (32)
Recalling the definition of $\varepsilon$ (eq. 9) and using the result before that in eq. 8, we obtain the following bound for the case of ${\mathrm{y}}=1$:
$\mathbb{P}\Big{(}\langle{\mathbf{x}},\widehat{\bm{\theta}}_{0}\rangle\leq 0\Big{)}\leq\exp\Bigg{(}-\frac{n\mu^{2}}{4}(1-2p)^{2}\textup{cos}^{2}\big{(}\beta({\mathbf{x}})\big{)}\Bigg{)}.$ (33)
This finishes the proof of the result for the case of ${\mathrm{y}}=1$. The result for the other case, i.e. ${\mathrm{y}}=-1$, can be derived similarly as above by analyzing with $r>0$ instead. In particular, for the case of ${\mathrm{y}}=-1$, we get:
$\mathbb{P}\Big{(}\langle{\mathbf{x}},\widehat{\bm{\theta}}_{0}\rangle>0\Big{)}\leq\exp\Bigg{(}-\frac{n\mu^{2}}{4}(1-2p)^{2}\textup{cos}^{2}\big{(}\beta({\mathbf{x}})\big{)}\Bigg{)}.$ (34)
(b) Upper bound $\widehat{\alpha}({\mathbf{x}})$. From eq. 8 and eq. 9, recall that:
$\langle{\mathbf{x}},\widehat{\bm{\theta}}_{0}\rangle=\frac{\varepsilon}{n\sigma^{2}},\text{ where }\varepsilon=\sum_{k=1}^{n}\widehat{\mathrm{y}}_{k}\langle{\mathbf{x}}_{k},{\mathbf{x}}\rangle.$ (35)
Again, we will prove the result for the case of ${\mathrm{y}}=1$ (i.e., ${\mathrm{x}}^{(1)}>0$); the result for the other case can be derived similarly.
Note that:
$\varepsilon=\sum_{k=1}^{n}\widehat{\mathrm{y}}_{k}\mathrm{x}_{k}^{(1)}\mathrm{x}^{(1)}+\underbrace{\sum_{k=1}^{n}\widehat{\mathrm{y}}_{k}\sum_{l=2}^{d}\mathrm{x}_{k}^{(l)}\mathrm{x}^{(l)}}_{\tilde{\varepsilon}}.$ (36)
Conditioned on $\widehat{\mathrm{y}}_{k}$ and $\mathrm{x}_{k}^{(1)}$, $\tilde{\varepsilon}\sim\mathcal{N}\big{(}0,n\sigma^{2}\sum_{l=2}^{d}(\mathrm{x}^{(l)})^{2}\big{)}$ using the fact that $\mathrm{x}^{(1)},\mathrm{x}^{(2)},\ldots,\mathrm{x}^{(d)}$ are all independent of each other. But $\sum_{l=2}^{d}(\mathrm{x}^{(l)})^{2}=\|{\mathbf{x}}\|^{2}-(\mathrm{x}^{(1)})^{2}=\|{\mathbf{x}}\|^{2}\Big{(}1-\textup{cos}^{2}\big{(}\beta({\mathbf{x}})\big{)}\Big{)}=\|{\mathbf{x}}\|^{2}\textup{sin}^{2}\big{(}\beta({\mathbf{x}})\big{)}$. So in other words, conditioned on $\widehat{\mathrm{y}}_{k}$ and $\mathrm{x}_{k}^{(1)}$, $\tilde{\varepsilon}\sim\mathcal{N}\big{(}0,n\sigma^{2}\|{\mathbf{x}}\|^{2}\textup{sin}^{2}\big{(}\beta({\mathbf{x}})\big{)}\big{)}$. Also, recalling that ${\mathrm{y}}_{k}=\text{sign}(\mathrm{x}_{k}^{(1)})$ and ${\mathrm{x}}^{(1)}>0$, we have that $\widehat{\mathrm{y}}_{k}\mathrm{x}_{k}^{(1)}\mathrm{x}^{(1)}=\widehat{\mathrm{y}}_{k}{\mathrm{y}}_{k}\big{|}\mathrm{x}_{k}^{(1)}\big{|}\big{|}\mathrm{x}^{(1)}\big{|}$. Let $\delta_{k}:=\widehat{\mathrm{y}}_{k}{\mathrm{y}}_{k}$. Using eq. 1, we have:
$\delta_{k}=\begin{cases}1\quad\text{ w.p. }1-p,\\ -1\quad\text{ o/w.}\end{cases}$ (37)
Moreover, $\delta_{k}$ is independent of $\mathrm{x}_{k}^{(1)}$. Using of all this, we get:
$\mathbb{P}\big{(}\varepsilon\leq 0\big{)}$ $=\mathbb{P}\Big{(}\sum_{k=1}^{n}\delta_{k}\big{|}\mathrm{x}_{k}^{(1)}\big{|}\big{|}\mathrm{x}^{(1)}\big{|}+\varepsilon\leq 0\Big{)}$ (38)
$=\mathbb{E}_{\big{\{}\delta_{k},\mathrm{x}_{k}^{(1)}\big{\}}_{k=1}^{n}}\Bigg{[}\Phi\Bigg{(}-\frac{\sum_{k=1}^{n}\delta_{k}\big{|}\mathrm{x}_{k}^{(1)}\big{|}\big{|}\mathrm{x}^{(1)}\big{|}}{\sqrt{n}\sigma\|{\mathbf{x}}\|\big{|}\textup{sin}\big{(}\beta({\mathbf{x}})\big{)}\big{|}}\Bigg{)}\Bigg{]}$ (39)
$\geq\mathbb{E}_{\big{\{}\delta_{k},\mathrm{x}_{k}^{(1)}\big{\}}_{k=1}^{n}}\Bigg{[}\Phi\Bigg{(}-\frac{\big{|}\sum_{k=1}^{n}\delta_{k}\big{|}\mathrm{x}_{k}^{(1)}\big{|}\big{|}\times\big{|}\mathrm{x}^{(1)}\big{|}}{\sqrt{n}\sigma\|{\mathbf{x}}\|\big{|}\textup{sin}\big{(}\beta({\mathbf{x}})\big{)}\big{|}}\Bigg{)}\Bigg{]}.$ (40)
Using the convexity of $\Phi(.)$ on $\mathbb{R}_{\leq 0}$, we can further lower bound eq. 40 by:
$\mathbb{P}\big{(}\varepsilon\leq 0\big{)}$ $\geq\Phi\Bigg{(}-\frac{\mathbb{E}\Big{[}\big{|}\sum_{k=1}^{n}\delta_{k}\big{|}\mathrm{x}_{k}^{(1)}\big{|}\big{|}\Big{]}\times\big{|}\mathrm{x}^{(1)}\big{|}}{\sqrt{n}\sigma\|{\mathbf{x}}\|\big{|}\textup{sin}\big{(}\beta({\mathbf{x}})\big{)}\big{|}}\Bigg{)},$ (41)
using Jensen’s inequality. Using the result of Lemma 2 and the fact that $\big{|}\mathrm{x}^{(1)}\big{|}=\|{\mathbf{x}}\|\big{|}\textup{cos}\big{(}\beta({\mathbf{x}})\big{)}\big{|}$ above, we get:
$\mathbb{P}\big{(}\varepsilon\leq 0\big{)}$ $\geq\Phi\Big{(}-{\big{(}1+\sqrt{n}\mu(1-2p)\big{)}\big{|}\textup{cot}\big{(}\beta({\mathbf{x}})\big{)}\big{|}}\Big{)}.$ (42)
Recalling eq. 35 and using the fact that $\Phi(-z)=\Phi^{\text{c}}(z)$ for all $z\in\mathbb{R}$ (recall that $\Phi^{\text{c}}(.)$ is the complementary CDF of a standard normal variable), we obtain the following bound for the case of ${\mathrm{y}}=1$:
$\mathbb{P}\Big{(}\langle{\mathbf{x}},\widehat{\bm{\theta}}_{0}\rangle\leq 0\Big{)}\geq\Phi^{\text{c}}\Big{(}{\big{(}1+\sqrt{n}\mu(1-2p)\big{)}\big{|}\textup{cot}\big{(}\beta({\mathbf{x}})\big{)}\big{|}}\Big{)}.$ (43)
Using Formula 7.1.13 from Abramowitz and Stegun, (1968), we have for any $z>0$:
$\Phi^{\text{c}}(z)>\sqrt{\frac{2}{\pi}}\Bigg{(}\frac{e^{-\frac{z^{2}}{2}}}{z+\sqrt{z^{2}+4}}\Bigg{)}.$
Further, using the fact that $z+\sqrt{z^{2}+4}<2(z+1)$ for $z>0$ above, we get:
$\Phi^{\text{c}}(z)>\frac{1}{\sqrt{2\pi}}\Bigg{(}\frac{e^{-\frac{z^{2}}{2}}}{z+1}\Bigg{)}.$ (44)
Now if $z\leq 2$, $z+1\leq 3$. Using this in eq. 44, we get:
$\Phi^{\text{c}}(z)>\frac{1}{3\sqrt{2\pi}}\exp\Big{(}-\frac{z^{2}}{2}\Big{)}.$ (45)
If $z>2$, $z+1<e^{z}<e^{z^{2}/2}$. Using this in eq. 44, we get:
$\Phi^{\text{c}}(z)>\frac{1}{\sqrt{2\pi}}\exp(-{z^{2}}).$ (46)
Combining eq. 45 and eq. 46, we get the following simpler bound:
$\Phi^{\text{c}}(z)>\frac{1}{3\sqrt{2\pi}}\exp(-{z^{2}}).$ (47)
Using this in eq. 43, we get:
@@
$\mathbb{P}\Big{(}\langle{\mathbf{x}},\widehat{\bm{\theta}}_{0}\rangle\leq 0\Big{)}\geq\frac{1}{3\sqrt{2\pi}}{\exp\Big{(}-{\big{(}1+\sqrt{n}\mu(1-2p)\big{)}^{2}\textup{cot}^{2}\big{(}\beta({\mathbf{x}})\big{)}}\Big{)}}.$ (48)
This finishes the proof of the result for the case of ${\mathrm{y}}=1$. The result for the other case (i.e., ${\mathrm{y}}=-1$) can be derived similarly as above; in particular, here we get:
@@
$\mathbb{P}\Big{(}\langle{\mathbf{x}},\widehat{\bm{\theta}}_{0}\rangle>0\Big{)}\geq\frac{1}{3\sqrt{2\pi}}{\exp\Big{(}-{\big{(}1+\sqrt{n}\mu(1-2p)\big{)}^{2}\textup{cot}^{2}\big{(}\beta({\mathbf{x}})\big{)}}\Big{)}}.$ (49)
This finishes the proof. ∎
Lemma 1.
For $t\leq\frac{1}{2}$, it holds that:
$e^{t}-1\leq t+t^{2}.$
Proof.
Using the Taylor expansion of $e^{t}$, we get:
$e^{t}-1$ $=t+\frac{t^{2}}{2!}+\frac{t^{3}}{3!}+\frac{t^{4}}{4!}\ldots$ (50)
$\leq t+\frac{t^{2}}{2}\big{(}1+t+t^{2}+\ldots\big{)}$ (51)
$=t+\frac{t^{2}}{2(1-t)}.$ (52)
Since $t\leq\frac{1}{2}$, $2(1-t)\geq 1$. Using this above gives us the desired result. ∎
Lemma 2.
For $p\leq\frac{1}{2}$, we have:
$\mathbb{E}\Big{[}\big{|}\sum_{k=1}^{n}\delta_{k}\big{|}\mathrm{x}_{k}^{(1)}\big{|}\big{|}\Big{]}\leq\sqrt{n}\sigma+n(1-2p)\mu\sigma.$
Proof.
First, note that:
$\mathbb{E}\Big{[}\big{|}\sum_{k=1}^{n}\delta_{k}\big{|}\mathrm{x}_{k}^{(1)}\big{|}\big{|}\Big{]}\leq\sqrt{\mathbb{E}\Bigg{[}\Big{|}\sum_{k=1}^{n}\delta_{k}\big{|}\mathrm{x}_{k}^{(1)}\big{|}\Big{|}^{2}\Bigg{]}}.$ (53)
Now using independence, we have:
$\mathbb{E}\Bigg{[}\Big{|}\sum_{k=1}^{n}\delta_{k}\big{|}\mathrm{x}_{k}^{(1)}\big{|}\Big{|}^{2}\Bigg{]}\leq\sum_{k}\mathbb{E}\Big{[}\big{|}\mathrm{x}_{k}^{(1)}\big{|}^{2}\Big{]}+\sum_{k\neq l}\mathbb{E}\big{[}\delta_{k}\big{]}\mathbb{E}\big{[}\delta_{l}\big{]}\mathbb{E}\Big{[}\big{|}\mathrm{x}_{k}^{(1)}\big{|}\Big{]}\mathbb{E}\Big{[}\big{|}\mathrm{x}_{l}^{(1)}\big{|}\Big{]}.$ (54)
As per our setting, $\mathbb{E}\Big{[}\big{|}\mathrm{x}_{k}^{(1)}\big{|}^{2}\Big{]}=\sigma^{2}$, $\mathbb{E}\Big{[}\big{|}\mathrm{x}_{k}^{(1)}\big{|}\Big{]}=\mu\sigma$ and $\mathbb{E}\big{[}\delta_{k}\big{]}=1-2p$. Using all of this in eq. 54, we get:
$\mathbb{E}\Bigg{[}\Big{|}\sum_{k=1}^{n}\delta_{k}\big{|}\mathrm{x}_{k}^{(1)}\big{|}\Big{|}^{2}\Bigg{]}\leq n\sigma^{2}+n(n-1)(1-2p)^{2}\mu^{2}\sigma^{2}\leq n\sigma^{2}+n^{2}(1-2p)^{2}\mu^{2}\sigma^{2}.$ (55)
Using this in eq. 53, we get:
$\mathbb{E}\Big{[}\big{|}\sum_{k=1}^{n}\delta_{k}\big{|}\mathrm{x}_{k}^{(1)}\big{|}\big{|}\Big{]}\leq\Big{(}n\sigma^{2}+n^{2}(1-2p)^{2}\mu^{2}\sigma^{2}\Big{)}^{1/2}\leq\sqrt{n}\sigma+n(1-2p)\mu\sigma.$ (56)
This finishes the proof. ∎
Appendix C Proof of Theorem 2
Proof.
Just like Theorem 1, we can prove the result for $\bm{\theta}^{\ast}={\mathbf{e}}_{1}$ and =$\mathcal{B}=\{{\mathbf{e}}_{1},{\mathbf{e}}_{2},\ldots,{\mathbf{e}}_{d}\}$ without loss of generality. In that case, ${\mathrm{a}}^{(j)}={\mathrm{x}}^{(j)}=\langle{\mathbf{x}},{\mathbf{e}}_{j}\rangle$ for $j\in[d]$, the ground truth label of ${\mathbf{x}}$ is $\text{sign}({\mathrm{x}}^{(1)})$, $\big{|}\textup{cos}\big{(}\beta({\mathbf{x}})\big{)}\big{|}=\frac{|{\mathrm{x}}^{(1)}|}{\|{\mathbf{x}}\|}$ and $\big{|}\textup{cot}\big{(}\beta({\mathbf{x}})\big{)}\big{|}=\frac{|{\mathrm{x}}^{(1)}|}{\sqrt{\sum_{j=2}^{d}({\mathrm{x}}^{(j)})^{2}}}$. Also note that:
$\text{acc}(\widehat{\bm{\theta}}_{0})=\mathbb{P}_{{\mathbf{x}}\sim\mathcal{D},\mathcal{T}}\Big{(}\textup{sign}\big{(}\langle{\mathbf{x}},\widehat{\bm{\theta}}_{0}\rangle\big{)}=\textup{sign}\big{(}{\mathrm{x}}^{(1)}\big{)}\Big{)}=\mathbb{E}_{{\mathbf{x}}\sim\mathcal{D},\mathcal{T}}\Big{[}\mathbbm{1}\Big{(}\textup{sign}\big{(}\langle{\mathbf{x}},\widehat{\bm{\theta}}_{0}\rangle\big{)}=\textup{sign}\big{(}{\mathrm{x}}^{(1)}\big{)}\Big{)}\Big{]}.$
(a) Lower bound. For the above case, we have:
$\text{acc}(\widehat{\bm{\theta}}_{0})$ $=\mathbb{E}_{{\mathbf{x}}\sim\mathcal{D},\mathcal{T}}\Bigg{[}\mathbbm{1}\Big{(}\textup{sign}\big{(}\langle{\mathbf{x}},\widehat{\bm{\theta}}_{0}\rangle\big{)}=\textup{sign}\big{(}{\mathrm{x}}^{(1)}\big{)}\Big{)}\Bigg{]}$ (57)

### v2

@@ B.1 Upper bound $\alpha_{0}(\mathbf{x})$
$\mathbb{P}\left((\sum_{i\in[n]}\xi_{i}(1+u_{i}))\langle y\bm{x},{\bm{{\mu}}}\rangle+\langle\bm{x},\tilde{\bm{z}}\rangle\leq 0\;\bigg{|}\;\{\xi_{i},y_{i},u_{i}\}_{i\in[n]}\right)=\Phi\left(-\frac{(\sum_{i\in[n]}\xi_{i}(1+u_{i}))\langle y\bm{x},{\bm{{\mu}}}\rangle}{\sqrt{n}\left\|\bm{\Sigma}^{1/2}\bm{x}\right\|_{\ell_{2}}}\right)\,,$
where $\Phi(t)=\frac{1}{\sqrt{2\pi}}\int_{-\infty}^{t}e^{-s^{2}/2}{\rm d}s$ denotes the cdf of standard normal distribution.
Combining the previous two equations, we arrive at
@@
$\mathbb{P}({\rm sign}(\langle\bm{x},{\widehat{\bm{\theta}}}_{0}\rangle)\neq y)$ $=\operatorname{\mathbb{E}}\left[\Phi\left(-\frac{(\sum_{i\in[n]}\xi_{i}(1+u_{i}))\langle y\bm{x},{\bm{{\mu}}}\rangle}{\sqrt{n}\left\|\bm{\Sigma}^{1/2}\bm{x}\right\|_{\ell_{2}}}\right)\right]$ (B.1)
$\geq\operatorname{\mathbb{E}}\left[\Phi\left(-\frac{|\sum_{i\in[n]}\xi_{i}(1+u_{i})|\;|\langle\bm{x},{\bm{{\mu}}}\rangle|}{\sqrt{n}\left\|\bm{\Sigma}^{1/2}\bm{x}\right\|_{\ell_{2}}}\right)\right]$
$\geq\Phi\left(-\frac{\operatorname{\mathbb{E}}[|\sum_{i\in[n]}\xi_{i}(1+u_{i})|]\;|\langle\bm{x},{\bm{{\mu}}}\rangle|}{\sqrt{n}\left\|\bm{\Sigma}^{1/2}\bm{x}\right\|_{\ell_{2}}}\right)$
where the expectation is with respect to $\{\xi_{i},u_{i}\}_{i\in[n]}$ and the last step follows by Jensen’s inequality and the fact that $\Phi(t)$ is a convex function for $t\leq 0$.
We next note that by Cauchy–Schwarz inequality,
$\operatorname{\mathbb{E}}\left[\Big{|}\sum_{i}\xi_{i}(1+u_{i})\Big{|}\right]$ $\leq\sqrt{\operatorname{\mathbb{E}}[(\sum_{i}\xi_{i}(1+u_{i}))^{2}]}$
$=\sqrt{\sum_{i,j}\operatorname{\mathbb{E}}[\xi_{i}\xi_{j}(1+u_{i})(1+u_{j})]}$
$=\sqrt{\sum_{i}\operatorname{\mathbb{E}}[(1+u_{i})^{2}]+\sum_{i\neq j}\operatorname{\mathbb{E}}[\xi_{i}]\operatorname{\mathbb{E}}[\xi_{j}]\operatorname{\mathbb{E}}[1+u_{i}]\operatorname{\mathbb{E}}[1+u_{j}]}$
$=\sqrt{5n+4n(n-1)(1-2p)^{2}}\leq\sqrt{5n}+\sqrt{5}n(1-2p)\,,$
where we used that $\xi_{i},\xi_{j}$ are independent for $i\neq j$, and independent from $u_{i}$’s. In addition, $\operatorname{\mathbb{E}}[\xi_{i}]=\operatorname{\mathbb{E}}[y_{i}\widehat{y}_{i}]=\mathbb{P}(y_{i}=\widehat{y}_{i})-\mathbb{P}(y_{i}\neq\widehat{y}_{i})=1-2p$. Also $\operatorname{\mathbb{E}}[u_{i}]\leq 1$ and $\operatorname{\mathbb{E}}[u_{i}^{2}]\leq 2$, since $u_{i}$ has unit sub-gaussian norm. Hence, we get
@@
$\mathbb{P}({\rm sign}(\langle\bm{x},{\widehat{\bm{\theta}}}_{0}\rangle)\neq y)\geq\Phi\left(-\frac{\sqrt{5}(1+\sqrt{n}(1-2p))\;|\langle\bm{x},{\bm{{\mu}}}\rangle|}{\left\|\bm{\Sigma}^{1/2}\bm{x}\right\|_{\ell_{2}}}\right)\,.$ (B.2)
We next use a classical lower bound on $\Phi$. Let $\Phi^{c}(t)=1-\Phi(t)=\Phi(-t)$. Using (Abramowitz and Stegun,, 1968, Equation (7.1.13)), we have for any $t>0$:
$\Phi^{\text{c}}(t)>\sqrt{\frac{2}{\pi}}\Bigg{(}\frac{e^{-\frac{t^{2}}{2}}}{t+\sqrt{t^{2}+4}}\Bigg{)}.$
Since $t+\sqrt{t^{2}+4}<2(t+1)$ for $t>0$ , we get:
$\Phi^{\text{c}}(t)>\frac{1}{\sqrt{2\pi}}\Bigg{(}\frac{e^{-\frac{t^{2}}{2}}}{t+1}\Bigg{)}>\frac{1}{2\sqrt{2\pi}}\exp(-{t^{2}}),$ (B.3)
where we used that $t+1\leq t^{2}+2\leq 2e^{t^{2}/2}$. Combining (B.3) and (B.2) we get
$\mathbb{P}({\rm sign}(\langle\bm{x},{\widehat{\bm{\theta}}}_{0}\rangle)\neq y)\geq\frac{1}{2\sqrt{2\pi}}\exp\left(-\frac{5(1+\sqrt{n}(1-2p))^{2}\langle\bm{x},{\bm{{\mu}}}\rangle^{2}}{\left\|\bm{\Sigma}^{1/2}\bm{x}\right\|_{\ell_{2}}^{2}}\right)\,,$
which completes the derivation of $\alpha_{0}(\bm{x})$.
B.2 Lower bound $\widetilde{\alpha}_{0}(\mathbf{x})$
We continue from (B.1), which reads
$\mathbb{P}({\rm sign}(\langle\bm{x},{\widehat{\bm{\theta}}}_{0}\rangle)\neq y)=\operatorname{\mathbb{E}}\left[\Phi\left(-\frac{(\sum_{i\in[n]}\xi_{i}(1+u_{i}))\langle y\bm{x},{\bm{{\mu}}}\rangle}{\sqrt{n}\left\|\bm{\Sigma}^{1/2}\bm{x}\right\|_{\ell_{2}}}\right)\right]\,,$ (B.4)
where the expectation is with respect to the randomness in training data, namely $u_{i}$ and $\xi_{i}$ for $i\in[n]$, while the test point $(\bm{x},y)$ is fixed.
Note that under our data model, $\langle y\bm{x},{\bm{{\mu}}}\rangle=(1+u)\left\|{\bm{{\mu}}}\right\|_{\ell_{2}}^{2}=(1+u)\gamma^{2}>0$ since $u$ is a non-negative random variable. We next define the following event:
$\mathcal{E}_{0}:=\left\{\sum_{i\in[n]}\xi_{i}(1+u_{i})\geq n(1-2p)/2\right\}\,.$ (B.5)
Recall that $\operatorname{\mathbb{E}}[\xi_{i}]=1-2p$ and by the Hoeffding-type inequality for sub-gaussian random variables,
$\mathbb{P}\left(\Big{|}\sum_{i}\xi_{i}(1+u_{i})-n(1-2p)(1+\operatorname{\mathbb{E}}[u_{i}])\Big{|}\geq t\right)\leq 2\exp\left(-\frac{t^{2}}{8n}\right)\,,$
(Note that $\|\xi_{i}(1+u_{i})\|_{\psi_{2}}\leq 1+\|u_{i}\|_{\psi_{2}}=2$.) By choosing $t=n(1-2p)/2$, and using the fact that $u_{i}>0$ for all $i$ we get
$\mathbb{P}\left(\sum_{i}\xi_{i}(1+u_{i})<n(1-2p)/2\right)\leq 2\exp\left(-\frac{n}{32}(1-2p)^{2}\right)\,.$
Hence, $\mathbb{P}(\mathcal{E}_{0}^{c})\leq 2\exp\left(-\frac{n}{32}(1-2p)^{2}\right)$.
We then have
$\operatorname{\mathbb{E}}\left[\Phi\left(-\frac{(\sum_{i\in[n]}\xi_{i}(1+u_{i}))\langle y\bm{x},{\bm{{\mu}}}\rangle}{\sqrt{n}\left\|\bm{\Sigma}^{1/2}\bm{x}\right\|_{\ell_{2}}}\right)\right]$ $=\operatorname{\mathbb{E}}\left[\Phi\left(-\frac{(\sum_{i\in[n]}\xi_{i}(1+u_{i}))\langle y\bm{x},{\bm{{\mu}}}\rangle}{\sqrt{n}\left\|\bm{\Sigma}^{1/2}\bm{x}\right\|_{\ell_{2}}}\right)(\mathbb{1}_{\mathcal{E}_{0}}+\mathbb{1}_{\mathcal{E}_{0}^{c}})\right]$
$\leq\operatorname{\mathbb{E}}\left[\Phi\left(-\frac{n(1-2p)\langle y\bm{x},{\bm{{\mu}}}\rangle}{2\sqrt{n}\left\|\bm{\Sigma}^{1/2}\bm{x}\right\|_{\ell_{2}}}\right)\mathbb{1}_{\mathcal{E}_{0}}\right]+\mathbb{P}(\mathcal{E}_{0}^{c})$
$=\Phi\left(-\frac{n(1-2p)\langle y\bm{x},{\bm{{\mu}}}\rangle}{2\sqrt{n}\left\|\bm{\Sigma}^{1/2}\bm{x}\right\|_{\ell_{2}}}\right)\mathbb{P}(\mathcal{E}_{0})+\mathbb{P}(\mathcal{E}_{0}^{c})$
$\leq\Phi\left(-\frac{n(1-2p)\langle y\bm{x},{\bm{{\mu}}}\rangle}{2\sqrt{n}\left\|\bm{\Sigma}^{1/2}\bm{x}\right\|_{\ell_{2}}}\right)+\mathbb{P}(\mathcal{E}_{0}^{c})\,,$
where in the first inequality we used the observation that $\langle y\bm{x},{\bm{{\mu}}}\rangle\geq 0$ as explained above, the definition of $\mathcal{E}_{0}$ and that $\Phi(-z)$ is a decreasing function in $z$.
We next recall the tail bound of normal distribution (Abramowitz and Stegun,, 1968, Equation (7.1.13)):
$\Phi^{c}(t)\leq\sqrt{\frac{2}{\pi}}\frac{e^{-t^{2}/2}}{t+\sqrt{t^{2}+\frac{8}{\pi}}}\leq\frac{1}{2}e^{-t^{2}/2}\,,$
for all $t\geq 0$, where $\Phi^{c}(t)=1-\Phi(t)=\Phi(-t)$ is the complementary CDF of normal variables. Using this bound and the bound we derived on $\mathbb{P}(\mathcal{E}_{0}^{c})$ we get
$\operatorname{\mathbb{E}}\left[\Phi\left(-\frac{(\sum_{i\in[n]}\xi_{i}(1+u_{i}))\langle y\bm{x},{\bm{{\mu}}}\rangle}{\sqrt{n}\left\|\bm{\Sigma}^{1/2}\bm{x}\right\|_{\ell_{2}}}\right)\right]\leq\frac{1}{2}\exp\left(-\frac{n(1-2p)^{2}\langle y\bm{x},{\bm{{\mu}}}\rangle^{2}}{8\left\|\bm{\Sigma}^{1/2}\bm{x}\right\|_{\ell_{2}}^{2}}\right)+2\exp\left(-\frac{n}{32}(1-2p)^{2}\right)\,.$
By invoking (B.4) we obtain